trey1v:

The temperature falls from 0 degrees to Negative 12 and one-fourth degrees in 3 and one-half hours. Which expression finds the change in temperature per hour?

Limitless2:

the answer your looking for is -12.25 / 3.5

trey1v:

thanks

somnium:

no its more than that

somnium:

its a linear expression

trey1v:

oh

somnium:

hold on js a sec

trey1v:

k

somnium:

so hes right about that part but thats your first step

somnium:

if you can solve for the piece of that, you get -3.5 ofc

somnium:

which is the change per hour, but you need a full expression

somnium:

so you would write something like this \[y=-3.5x\]

trey1v:

hm

somnium:

let x be the number of hours, and y is the tempurature after x hours

trey1v:

im on edgenuity..

somnium:

whats the platform change?

trey1v:

1 attachment
somnium:

ah then your answer is d

trey1v:

oh

somnium:

you just have to convert the given improper fractions into proper fractions, and compare

trey1v:

mk

somnium:

and ill assume you know how to convert them?

somnium:

if not i can explain it

trey1v:

yh

ykbatman:

-12.25

ykbatman:

is yo answer

MysticEcho:

-12.25 is the answer bro!!!!

adrianluvvsyouu2:

Change in temperature is 0 - 12 and 1/4, therefore just -12 and 1/4 which is equal to -49/4 Then if time is 3 and 1/2 hours, that fraction is the same as 1/2 + 6/2 = 7/2 hours (-49/4)/(7/2) = the answer, but we need to simplify. When dividing by any fraction, it is equivalent as multiplying by its reciprocal (The flipped version) so 7/2 becomes *2/7 -49/4 * 2/7 = -98/28 Both the top and bottom can be divided by 14 to simplify it to -7/2

adrianluvvsyouu2:

The reason why it's not 12.25 (or 12 and 1/4) is because we have to find it PER hour. Rate is always a unit that is divided by time.

grimreaper124:

lol